### v3

# Language-guided Scale-aware MedSegmentor for Lesion Segmentation in Medical Imaging

Abstract
In clinical practice, segmenting specific lesions based on the needs of physicians can significantly enhance diagnostic accuracy and treatment efficiency. However, conventional lesion segmentation models lack the flexibility to distinguish lesions according to specific requirements. Given the practical advantages of using text as guidance, we propose a novel model, Language-guided Scale-aware MedSegmentor (LSMS), which segments target lesions in medical images based on given textual expressions. We define this as a new task termed Referring Lesion Segmentation (RLS). To address the lack of suitable benchmarks for RLS, we construct a vision-language medical dataset named Reference Hepatic Lesion Segmentation (RefHL-Seg). LSMS incorporates two key designs: (i) Scale-Aware Vision-Language attention module, which performs visual feature extraction and vision-language alignment in parallel. By leveraging diverse convolutional kernels, this module acquires rich visual representations and interacts closely with linguistic features, thereby enhancing the model’s capacity for precise object localization. (ii) Full-Scale Decoder, which globally models multi-modal features across multiple scales and captures complementary information between them to accurately delineate lesion boundaries. Additionally, we design a specialized loss function comprising both segmentation loss and vision-language contrastive loss to better optimize cross-modal learning. We validate the performance of LSMS on RLS as well as on conventional lesion segmentation tasks across multiple datasets. Our LSMS consistently achieves superior performance with significantly lower computational cost. Code and datasets will be released.
Index Terms:
vision-language, medical image, segmentation, multi-scale
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Figure 1: Comparison between Referring Lesion Segmentation and conventional lesion segmentation tasks. In the images displaying segmentation results, the regions highlighted in red represent the Ground Truth. For intuitive correspondence with the left-right references in the text, all CT images in this paper have been mirrored horizontally.
The significance of lesion segmentation in medical image analysis has been widely recognized [1, 2]. It enables the precise identification and delineation of pathological regions, which is essential for accurate diagnosis and treatment planning. As shown in Fig. 1(a), the Classical Lesion Segmentation task involves inputting a medical image and obtaining segmentation results encompassing lesions within the image [3, 4]. With the advancement of multi-modal research, studies have emerged that incorporates textual input as supplementary information for segmentation. Fig. 1(b) illustrates Text-Augmented Lesion Segmentation, wherein diagnostic texts provided by physicians aid in image interpretation [5]. This task involves inputting a medical image along with its diagnostic text and outputting segmentation results for All Lesions without distinction. However, in clinical practice, physicians often need to segment specific lesions to provide tailored diagnosis and treatment, thus rendering Conventional Lesion Segmentation  [^1] In this paper, Conventional Lesion Segmentation is used as a collective term to refer to the two tasks illustrated in Figure 1: (a) Classical Lesion Segmentation and (b) Text-Augmented Lesion Segmentation.
tasks inadequate for practical applications. For instance, in radiation therapy, the accuracy of tumor segmentation directly influences the precision of radiation beam targeting, which determines the effectiveness of the treatment and the potential risk of side effects. As shown in Fig. 1(a) and (b), Conventional Lesion Segmentation tasks treat all lesions equally, which is insufficient for clinical scenarios. Text, as a convenient medium for expressing physicians’ needs, can be used to indicate target segmentation objects. Therefore, we introduce a new task of Referring Lesion Segmentation (RLS), where medical images are accompanied by a language expression that indicate a Specific Lesion within the image. Fig. 1(c) represents the introduced task, RLS, which involves segmenting the largest lesion in the right liver as described by the language expression. Compared to the task in Fig. 1(b), RLS emphasizes the ability to differentiate lesions based on the text. To support evaluation of RLS, we developed Reference Hepatic Lesion Segmentation (RefHL-Seg) dataset—the first benchmark specifically designed for this new task.
Figure 2: Comparison of Transformer-based architectures. (a) Existing architectures for related tasks. (b) our LSMS for RLS.
Deep learning methods have demonstrated outstanding performance in medical image segmentation tasks. Classical Lesion Segmentation methods often relies on architectures such as U-Net [1] or Transformer [6]. Methods combining these two architectures generally down-sample feature maps to acquire high-level contextual information, followed by up-sampling to reconstruct spatial dimensions [3, 7]. To further enhance segmentation performance, approaches incorporating language modality have emerged [5, 8]. Recently, Transformer-based models have shown great promise in this area, as their ability to model long-range dependencies facilitates effective integration of visual and linguistic information. As shown in Fig. 2(a), existing Transformer-based methods include a fusion module at the end of each stage to integrate the extracted visual features with linguistic information. Models utilizing this architecture include the mature Natural Image Referring Segmentation model LAVT [9] and the recent Text-Augmented Lesion Segmentation model LViT [5]. By introducing a multi-scale structure, they effectively exploit rich vision-language knowledge. However, these methods are not directly applicable to the RLS task. Existing methods treat visual feature extraction and cross-modal fusion as two independent steps, leaving room for improvement in achieving visual-linguistic alignment within the semantic space. During decoding, they adopt sequential structures that tend to produce single-scale representations at each level, while enhancing inter-scale interaction may help capture core semantic information.
Upon analyzing the requirements of RLS and previous efforts in related tasks, we identify two key aspects that address the unique challenges of RLS while aligning with conventional segmentation goals: (i) Robust vision-language modeling. In the medical visual environment, the notable variations in size and shape among objects make it crucial to effectively model visual-linguistic consistency for accurate object localization. Fusing visual and linguistic features after visual feature extraction may overlook the rich local visual information correlated with linguistic guidance, thereby impacting the model’s object localization performance. (ii) Comprehensive multi-scale interaction. Globally modeling the complex differences across scales enables the extraction of optimal global visual-linguistic features. Given the complexity of the visual environment in medical images, neglecting complementary information between scales may result in insufficient capability to identify lesion boundaries during segmentation.
In light of the aforementioned analysis, we propose a model named Language-guided Scale-aware MedSegmentor (LSMS), as illustrated in Fig. 2(b). Within LSMS, we introduce a Scale-aware Vision-Language Attention (SVLA) module embedded in the encoder block. SVLA captures scale-aware visual knowledge and models visual-linguistic relationships in an integrated manner, enhancing the visual-linguistic alignment in the semantic space. As shown in Fig. 3, LSMS (w/o FSD) equipped with SVLA exhibits a significant enhancement in lesion localization capability compared to the results of existing vision-language model LViT [5] and LAVT [9]. Additionally, we devise a Full-Scale Decoder (FSD) that globally models multi-modal information by aligning and integrating multi-modal feature maps from various scales, thereby enhancing the comprehension of details within complex medical images. In Fig. 3(b), LSMS, when contrasted with LSMS (w/o FSD), exhibits a more precise prediction of lesion boundaries during segmentation.
Figure 3: Qualitative results of different approaches. The red regions denote the Ground Truth, while the green regions represent the segmentation results of our LSMS, LSMS (w/o FSD), LViT [5] and LAVT [9]. In sample (b), for ease of observation, the key region within the image have been enlarged, with a yellow rectangular box serving as a reference for location.
Additionally, we have designed a specialized loss function to constrain the model’s training, which includes the Segmentation Loss $\mathcal{L}_{Seg}$ for the segmentation results and the Vision-Language Contrastive Loss $\mathcal{L}_{Con}$ for the linguistic features and final multi-modal features. Specifically, $\mathcal{L}_{Seg}$ enhances the focus on the object boundaries, thereby improving edge detection capabilities. Meanwhile, $\mathcal{L}_{Con}$ further aligns the visual knowledge of the target region with the linguistic information, leading to improved accuracy in target localization.
In summary, our contributions encompass four aspects:
  1. 1.
    We introduce a new task of RLS, which entails segmenting the target object in medical images based on the language expression. We have established the RefHL-Seg dataset as a new benchmark for RLS.
  2. 2.
    We propose LSMS for lesion segmentation, comprising a SVLA module to improve object localization capability and a full-scale decoder to enhance the accuracy of lesion boundary prediction.
  3. 3.
    We design a specialized loss function to optimize fine-grained discrimination and visual-linguistic alignment, thereby enhancing the accuracy.
  4. 4.
    We conduct comprehensive experiments on the RefHL-Seg dataset for RLS, as well as on datasets for conventional lesion segmentation tasks. Experimental results demonstrate the superiority of LSMS over current state-of-the-art methods with lower computational costs.
The preliminary work was presented as a conference paper at the International Joint Conference on Artificial Intelligence (IJCAI) 2023 [10]. This paper extends the prior work by introducing a new task (RLS), a dataset (RefHL-Seg) as the new benchmark, as well as method and experimental enhancements. We optimize the scale-wise module design in the model and propose a specialized loss function, which improves the model’s performance in the domain of lesion segmentation. Furthermore, we apply the proposed model to the newly introduced RLS task and validate it on conventional lesion segmentation benchmarks, demonstrating the generalizability of our approach.
II Related Works
II-A Medical Image Segmentation
For medical image segmentation tasks, early methods [11, 12] extended from image classification networks to achieve semantic segmentation, among which the Fully Convolutional Network (FCN) [11] is an end-to-end semantic segmentation network [18]. Multi-scale architectures have been proven to enhance segmentation performance based on CNN methods. U-Net [1] is considered a pioneer in medical image segmentation. U-Net++ [2] further improved U-Net by enhancing skip connections to bridge the semantic gap between shallow encoder layers and deep decoder layers. MS-DualGuided [13] focused on both spatial and channel dimensions of feature maps at different scales. CA-Net [14] proposed a multi-scale context-aware network for multi-modal medical image segmentation. Subsequently, the introduction of Transformer [6] was found to enhance medical image segmentation performance. TransUNet [3] integrated the power of Transformers with the U-Net architecture for improved instance segmentation. Swin-Unet [7] specifically designed to leverage the hierarchical features and shifted window mechanisms of Swin Transformer [15], thereby enhancing the model’s capability to capture both local and global context. Furthermore, LViT [5] proposed a medical image segmentation model with textual assistance, introducing text input as auxiliary information during the encoding stage of segmentation. RecLMIS [8] proposes a novel cross-modal conditioned method for Text-Augmented Lesion Segmentation. While current research incorporates text as an aid for image understanding, existing benchmarks and methods are unable to distinguish specific objects in medical images, thereby falling short of fully addressing the practical needs of physicians.
Figure 4: (a) An illustration of LSMS. Initially, the input image and the reference expression are embedded separately through the vision embedding block and the BERT [16] language encoder, yielding visual feature $V_{1}$ and linguistic feature $L$, which are then fed into the language-guided vision encoder. The encoder incorporates the Scale-aware Vision-Language Attention (SVLA) module to interact between visual knowledge from different receptive fields and linguistic features. The encoder blocks at each stage learn rich multi-modal features $F_{i},i\in\{1,2,3,4\}$, which are subsequently fed into the full-scale decoder. Through the Position Alignment block, $F_{i},i\in\{1,2,3,4\}$ uniformly resize the feature maps of various scales while preserving channel disparities, resulting in $P_{i},i\in\{1,2,3,4\}$. Features $P_{i},i\in\{2,3,4\}$ are then globally modeled across scales for final segmentation. (b) An illustration of the operational mechanism of the vision-language contrastive loss $\mathcal{L}_{Con}$.
II-B Natural Image Referring Segmentation
Early referring segmentation methods in natural images [17, 18, 19] typically combined visual and linguistic features by concatenation, utilizing FCNs for cross-modal feature learning and prediction. In contrast, attention-based methods, such as vision-guided linguistic attention [20] and Cross-Modal Self-Attention [21], focus on aligning visual content with linguistic information. Bi-directional relationship networks [22] capture mutual guidance, while others [23, 24] use sentence structure to model cross-modal attributes. More recently, Transformer-based methods have enhanced the modeling of long-range cross-modal dependencies, leading to significant advancements in referring segmentation. VLT [25] and EFN [26] utilize a Transformer-based encoder-decoder framework, employing attention mechanisms in decoding stages to augment contextual information. LAVT [9] adopt an early fusion approach, modeling multi-modal context within the Transformer encoders. Prompt-RIS [27] leverages bidirectional prompting for better vision-language interaction. Current methods for Natural Image Referring Segmentation treat visual feature extraction and cross-modal fusion as two separate steps, and use sequential upsampling structures during decoding. However, due to the complexity of medical visual environments, these methods struggle to meet the requirements of RLS. In this paper, we propose the SVLA, which models visual information and visual-linguistic relationships in an integrated manner, and employ a full-scale decoder to promote comprehensive understanding of multi-scale knowledge.
III Language-Guided Scale-Aware Medical Segmentor
III-A Overview
In the proposed Language-guided Scale-aware MedSegmentor (LSMS), we learn visual knowledge from diverse receptive fields and tightly engage with linguistic features. Subsequently, leveraging a full-scale decoder, we globally model multi-modal information across all scales, thereby facilitating RLS task. The training loss includes the Segmentation Loss $\mathcal{L}_{{Seg}}$ and the vision-language contrastive loss $\mathcal{L}_{{Con}}$. The overall architecture of LSMS is presented in Fig. 4.
Given an image and a language expression, our model predicts the segmentation mask corresponding to the language reference. Our LSMS follows a workflow composed of [language-guided vision encoder] - [full-scale decoder]. LSMS comprises four stages, each with varying numbers of encoder blocks and different feature map resolutions. The encoder (Sec. III.B) incorporates a novel Scale-aware Vision-Language Attention module (Sec. III.C), which learns rich visual knowledge through convolutions of different sizes and closely interacts with linguistic features. Additionally, we propose a full-scale decoder (Sec. III.D) to globally model multi-modal information across multiple scales, enhancing the understanding of context details. During training, we employ the $\mathcal{L}_{{Seg}}$ to supervise the segmentation results and the $\mathcal{L}_{{Con}}$ to constrain the linguistic features and the final multi-modal representations (Sec. III.E). In the following subsections, we elaborate on each component of LSMS.
III-B Language-Guided Vision Encoder
To facilitate deep interaction between linguistic features and the complex visual environment of medical images, our encoder leverages a novel attention module (Sec. III.C) to perceive visual details in the image under linguistic guidance, thus obtaining valuable multi-modal representations. The structure of the encoder block is depicted in Fig. 5(a).
As illustrated in Fig. 4, our encoding phase comprises four stages with decreasing feature map resolutions. The $i$-th stage consists of $N_{i}$ encoder blocks. For the sake of clarity, we assume that $N_{i}=1,i\in\{1,2,3,4\}$ in this section. LSMS receives an input image alongside a reference expression. We extract the linguistic feature $L\in\mathbb{R}^{C_{l}\times T}$ using the BERT [16] language encoder, where $T$ denotes the number of words, and $C_{l}$ represents the channel number of the linguistic feature. Similarly, the input image undergoes an embedding block to yield the initial visual input $V_{1}\in\mathbb{R}^{C_{v1}\times H_{1}\times W_{1}}$ for the encoder, where $H_{1}$ and $W_{1}$ represent the height and width of the visual feature, and $C_{v1}$ represents the channel number. Each stage comprises a down-sampling block and a stack of encoder blocks. For each stage, the aggregation of the multi-modal feature maps $F_{i}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ can be expressed as:
[…]
As depicted in Fig. 5(a), the architecture of encoder blocks follows the design of ViT [28], but we introduce a SVLA module (Sec. III.C) to replace the conventional self-attention mechanism. The workflow of the encoder block is illustrated by the following:
[…]
TABLE I: Detailed settings of different stages in our LSMS.
Stage 1 Stage 2 Stage 3 Stage 4
number of blocks ($N_{i}$) 3 3 5 2
visual output size $\frac{H}{4}\times\frac{W}{4}$ $\frac{H}{8}\times\frac{W}{8}$ $\frac{H}{16}\times\frac{W}{16}$ $\frac{H}{32}\times\frac{W}{32}$
channel 64 128 320 512
III-C Scale-aware Vision-Language Attention
In medical images, instances vary greatly in size and shape, posing a challenge in pinpointing lesions referred to linguistic cues within intricate visual contexts. Addressing this, we learn scale-aware visual knowledge from diverse receptive fields by employing convolutional kernels of varying sizes, integrating linguistic features with visual knowledge across multiple scales.
As depicted in Fig. 5(b), our proposed attention mechanism, termed Scale-aware Vision-Language Attention (SVLA), initially captures preliminary visual features through a convolution operation, and then employs Visual Knowledge Branch and Language-Guided Branch to model rich visual knowledge and visual-linguistic relationships, followed by a $1\times 1$ convolution operation to learn the interplay between the branches. In $i$-th stage, with visual input $V_{i}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ and language input $L\in\mathbb{R}^{C_{l}\times T}$, we obtain the preliminary visual feature map $V_{i}^{pre}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ by the formula $V_{i}^{pre}=Conv_{5\times 5}(V_{i})$.
Visual Knowledge Branch
To accommodate the characteristics of medical image visual environment, we devised ConvUnits to capture scale-aware visual knowledge. ConvUnits comprises diverse convolution kernels, with each unit consisting of a $1\times d_{j}$ and a $d_{j}\times 1$ convolution operations, where $j\in\{1,2,3\}$. Each SVLA module comprises three ConvUnits, aimed at capturing scale-aware visual knowledge from various receptive fields. Visual knowledge $Att_{i}^{V}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ can be derived using the following formula:
$Att_{i}^{V}=\sum_{j=1}^{3}ConvUnit_{j}(V_{i}^{pre}),$ (4)
where $ConvUnit_{j}(\cdot)$ indicates $j$-th ConvUnit. The rectangular convolution kernels in ConvUnits enable the acquisition of detailed visual information with low computational costs.
Figure 5: (a) An illustration of the encoder block in the Language-Guided Vision Encoder. (b) An illustration of the Scale-aware Vision-Language Attention module.
Language-Guided Branch
We employ the Language-Guided Branch to model relationships between linguistic information and various visual coordinates, facilitating the guidance of lesion localization in the complex visual environment. The steps to obtain language-guided knowledge $Att_{i}^{L}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ are as follows:
$V_{i1},V_{i2}=flatten(\omega_{v1}(V_{i}^{pre})),\omega_{v2}(V_{i}^{pre}),$ (5)
$L_{i1},L_{i2}=\omega_{l1}(L),\omega_{l2}(L),$ (6)
$\alpha_{i}=V_{i1}^{\top}L_{i1},$ (7)
$Att_{i}^{L\prime}=unflatten((softmax(\frac{\alpha_{i}}{\sqrt{C_{l}}})L_{i2}^{\top})^{\top}),$ (8)
$Att_{i}^{L}=\omega_{f}(Att_{i}^{L\prime})\odot V_{i2},$ (9)
where $\omega_{v1}$, $\omega_{v2}$, $\omega_{l1}$, $\omega_{l2}$, $\omega_{f}$ are projection functions, $flatten(\cdot)$ denotes the operation of flattening the two spatial dimensions into a single dimension along the rows, $unflatten(\cdot)$ indicates the opposite operation of $flatten(\cdot)$, and $\odot$ is element-wise matrix multiplication operation. $\omega_{l1}$ and $\omega_{l2}$ each is implemented as a 1×1 convolution, yielding channels of size $C_{vi}$. $\omega_{v1}$, $\omega_{v2}$ and $\omega_{f}$ each is defined as a 1×1 convolution and an instance normalization.
Comprehensive Attention
We integrate information from the Visual Knowledge Branch and the Language-Guided Branch to compute comprehensive attention weights through convolution, thereby reweighting the input $V_{i}$ to the SVLA module. The feature $F_{i}^{Att}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ is obtained using the following:
$F_{i}^{Att}=Conv_{1\times 1}(V_{i}^{pre}+Att_{i}^{V}+Att_{i}^{L})\odot V_{i},$ (10)
where $Conv_{1\times 1}$ represents the $1\times 1$ convolution operation, and $\odot$ is element-wise matrix multiplication operation.
III-D Full-Scale Decoder
To accommodate the complexity of medical visual environment, a comprehensive cross-scale understanding of visual-linguistic contexts is necessary. Therefore, we devised a Full-Scale Decoder (FSD) to capture high-level semantics following the encoder. Unlike previous methods [29, 30, 31] with sequential structures, we globally process multi-modal features of various scales after aligning their positions. We employ Position Alignment to map them to the same feature map size of $F_{1}$ while retaining their original channel numbers. Subsequently, we concatenate features from different scales, pass them through a lightweight InterScale block for globally modeling multi-scale contexts, and finally generate segmentation predictions. The process is as follows:
$P_{1},P_{2},P_{3},P_{4}=PositionAlign(F_{1},F_{2},F_{3},F_{4}),$ (11)
$Out=Seg(InterScale(Concat[P_{2},P_{3},P_{4}])),$ (12)
where $PositionAlign(\cdot)$ represents the Position Alignment operation, $InterScale(\cdot)$ is implemented as a lightweight Hamburger [32] function, and $Seg(\cdot)$ indicates a $1\times 1$ convolution and an up-sampling function for final prediction. $PositionAlign(\cdot)$ is realized through bilinear interpolation operations.
It is noteworthy that we exclusively utilize features generated from $2,3,4$-th stages for global decoding. This choice is informed by the observation that the shallow features from the first stage exhibit a lower degree of visual-linguistic consistency, encompassing redundant information from irrelevant regions in the images. This redundancy hampers lesion localization guided by linguistic cues and interferes with the segmentation of target lesions in the visual environment. The efficacy of this design will be validated in the Ablation Study (Sec. IV.E). Further visualization analysis (Sec. IV.F) will be conducted to explore the disparities and characteristics of features from different stages.
III-E Loss Function
We design a specialized loss function to constrain the model’s training, which includes the Segmentation Loss $\mathcal{L}_{Seg}$ for the segmentation results and the Vision-Language Contrastive Loss $\mathcal{L}_{Con}$ for the linguistic features and final multi-modal features.
Specifically, we perform dilation and erosion operations on the ground truth label $M^{gt}$ to obtain the boundary region $E$, defined as $E=dilate(M^{gt})-erode(M^{gt})$. Here, $dilate(\cdot)$ expands and $erode(\cdot)$ contracts the object’s boundaries. A weight $\lambda$ is then assigned to the edge region where $E^{i}=1$. The loss function for the segmentation mask is defined as:
$\mathcal{L}_{Seg}=\mathcal{L}_{focal}(M,M^{gt})+W\odot\mathcal{L}_{dice}(M,M^{gt}),$ (13)
where $W^{i}=\lambda E^{i}+(1-E^{i})$, $\mathcal{L}_{focal}$ and $\mathcal{L}_{dice}$ are focal loss [33] and dice loss [34].
Furthermore, $\mathcal{L}_{Con}$ aligns the visual knowledge of the target region with the linguistic information while simultaneously repelling the visual representations of the target and non-target regions, as illustrated in Figure 3(b). In the figure, $S_{align}$ measures the similarity between the features of positive pixels and the average linguistic feature, while $S_{repel}$ quantifies the dissimilarity between the features of positive and negative pixels. The formula for $\mathcal{L}_{Con}$ is as follows:
$\mathcal{L}_{Con}$ $=-\frac{1}{|\mathcal{P}|}\sum_{i}^{|\mathcal{P}|}\frac{e^{(p_{i}\cdot L_{avg}/\tau)}}{e^{(p_{i}\cdot L_{avg}/\tau)}+\sum_{j}^{|\mathcal{N}|}e^{(p_{i}\cdot n_{j}/\tau)}},$ (14)
where $p_{i}$ represents the feature of the $i$-th positive pixel in the positive pixel set $\mathcal{P}$, $n_{j}$ denotes the feature of the $j$-th negative pixel in the negative pixel set $\mathcal{N}$, and $\tau$ is a hyperparameter that controls the sharpness of the probability distribution. $L_{avg}$ denotes the average pooled linguistic feature, computed as $L_{avg}={proj}\left(\frac{1}{T}\sum^{T}L_{t}\right)$. $L_{t}$ is the $t$-th linguistic token.
Our joint training loss is defined as follows:
$\mathcal{L}=\alpha\mathcal{L}_{Con}+\mathcal{L}_{Seg},$ (15)
where $\alpha$ is a hyperparameter.
TABLE II: Experimental results on the RefHL-Seg, QaTa-COV19 and MosMedData+ datasets in terms of Dice and mIoU. The best scores are in red, and the secondbest scores are in blue.
RefHL-Seg QaTa-COV19 MosMedData+
Method Backbone Text Param (M) Flops (G) Dice (%) mIoU (%) Dice (%) mIoU (%) Dice (%) mIoU (%)
U-Net [1] CNN 14.8 50.3 48.67 37.89 79.02 69.46 64.60 50.73
UNet++ [2] CNN 74.5 94.6 51.84 40.27 79.62 70.25 71.75 58.39
AttUNet [35] CNN 34.9 101.9 50.67 40.11 79.31 70.04 66.34 52.82
nnUNet [36] CNN 19.1 412.7 51.63 41.89 80.42 70.81 72.59 60.36
TransUNet [3] Hybrid 105 56.7 52.33 44.76 78.63 69.13 71.24 58.44
Swin-Unet [7] Hybrid 82.3 67.3 51.34 44.26 78.07 68.34 63.29 50.19
UCTransNet [4] Hybrid 65.6 63.2 54.87 44.38 79.15 69.60 65.90 52.69
LViT [5] Hybrid 28.0 54.0 54.84 44.65 81.12 71.37 72.58 60.40
LSMS Hybrid 8.78 8.91 58.75 49.39 82.14 72.07 73.14 61.76
ConVIRT [37] CNN 35.2 44.6 61.37 53.56 79.72 70.58 72.06 59.73
TGANet [38] CNN 19.8 41.9 63.28 55.49 79.87 70.75 71.81 59.28
GLoRIA [39] Hybrid 45.6 60.8 63.69 55.82 79.94 70.68 72.42 60.18
ViLT [40] Hybrid 87.4 55.9 64.58 56.65 79.63 70.12 72.36 60.15
CLIP [41] Hybrid 87.0 105.3 68.76 60.33 79.81 70.66 71.97 59.64
LAVT [9] Hybrid 118.6 83.8 69.03 60.98 79.28 69.89 73.29 60.41
SLViT [10] Hybrid 103.4 50.1 71.94 62.83 80.68 71.57 73.08 60.11
Prompt-RIS [27] Hybrid 225.9 108.7 72.11 63.08 81.13 71.94 74.12 60.97
LViT [5] Hybrid 29.7 54.1 71.48 62.02 83.66 75.11 74.57 61.33
RecLMIS [8] CNN 23.7 24.1 73.01 63.59 85.22 77.00 77.48 65.07
LSMS (1/4) Hybrid 25% 8.78 8.91 77.63 69.02 84.97 76.98 77.46 66.02
LSMS (1/2) Hybrid 50% 8.78 8.91 78.33 69.76 85.21 77.08 77.65 66.24
LSMS Hybrid 100% 8.78 8.91 78.24 70.31 85.57 77.60 78.32 67.41
IV Experiments
IV-A Datasets
We conducted experiments on three datasets to assess the performance of LSMS: our self-established dataset, Reference Hepatic Lesion Segmentation (RefHL-Seg), as well as MosMedData+ [42, 43] and QaTa-COV19 [44] datasets.
RLS, as a new task, lacks suitable benchmarks. The benchmarks established for Natural Image Referring Segmentation cannot be directly transferred to the medical domain due to fundamental differences in acquisition conditions and visual properties. Medical images differ from natural images in texture, contrast, and annotation complexity, requiring expert knowledge. Their targets often have unclear boundaries and varying sizes, necessitating multi-scale processing. To tackle these issues, we developed the Reference Hepatic Lesion Segmentation (RefHL-Seg) dataset, specifically designed for the RLS task.
As the first dataset tailored specifically for RLS, RefHL-Seg comprises 2,283 abdominal CT slices from 231 cases, predominantly featuring lesions such as Hemangiomas, Liver Cysts, Hepatocellular Carcinomas (HCC), Focal Nodular Hyperplasia, and Metastasis. With collaboration from radiology experts, we meticulously annotated all liver lesions for the first time, providing detailed descriptions of their locations and morphologies. Each lesion’s textual annotation includes information such as liver segment (location), diameter, shape, boundary, enhancement pattern, lesion type, and more. Corresponding segmentation masks were delineated. An exemplary textual annotation containing all relevant information is provided as follows: “A vascular tumor in segment VI of the liver, with a diameter of 7.5mm, irregular shape, clear boundary, and non-ring enhancement.” Additionally, experiments will involve testing language expressions that contain only partial information sufficient for lesion localization, such as ”A vascular tumor in segment VI of the liver, with an irregular shape.”
The MosMedData+ [42, 43] and QaTa-COV19 [44] datasets are established for Classical Lesion Segmentation tasks. Recent study [5] extended the textual annotations of these datasets, serving for the evaluation of Text-Augmented Lesion Segmentation. The MosMedData+ dataset comprises 2,729 CT scan slices of lung infections, primarily containing information about the location of lung infections and the number of infected regions. The QaTa-COV19 dataset consists of 9,258 chest X-ray images manually annotated with COVID-19 lesions, focusing on whether both lungs are infected, the quantity of lesions, and the approximate location of the infected regions.
IV-B Evaluation Metric
In line with prior research [5], we utilize the Dice score and mean Intersection-over-Union (mIoU) to assess the effectiveness of our proposed method. The Dice score quantifies performance by calculating the intersection of the predicted results and the ground truth, divided by twice the sum of the sizes of the two sets. The mIoU computes the average Intersection over Union for multiple segmentation instances, providing an overall measure of segmentation accuracy across all instances.
TABLE III: Ablation studies on the RefHL-Seg val set. The optimal scores are highlighted in red.
Dice mIoU
(a) Kernel size of the single ConvUnit in SVLA
$d_{1}=5$ 72.95 64.14
$d_{1}=7$ 73.28 64.81
$d_{2}=11$ 74.03 65.27
$d_{2}=15$ 73.97 65.31
$d_{3}=19$ 72.33 63.97
$d_{3}=21$ 73.36 65.22
(b) Ablation on design choices of SVLA
ConvU1 ConvU2 ConvU3 PM
76.75 67.34
76.89 67.45
76.65 67.22
77.23 68.91
78.34 70.31
(c) Effectiveness of FSD
LSMS (w/o FSD) 74.92 65.77
LSMS (w/ FSD) 78.34 70.31
(d) Ablation on design choices of FSD
$P_{4}$-Head-MLP 76.03 67.58
$P_{1},P_{2},P_{3},P_{4}$-MLP-Concat-MLP 77.41 68.79
$P_{1},P_{2},P_{3},P_{4}$-Concat-Head-MLP 77.97 69.38
$P_{1},P_{2},P_{3},P_{4}$-MLP-Concat-Head-MLP 76.40 68.29
(e) FSD on various stages
$P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$
77.94 66.59
78.80 67.54
77.89 67.36
78.34 70.31
77.97 69.38
IV-C Implementation Details
The proposed LSMS is implemented using PyTorch, leveraging the BERT implementation from the HuggingFace Transformer library [45]. Regarding dataset partitioning, we separated the original training set into training and validation sets. For the convolutional layers in the Visual Knowledge Branch of SVLA, we initialized the weights using SegNeXt [46] pre-trained on ImageNet-22K [47]. The language encoder of LSMS was initialized with the official pre-trained BERT weights, consisting of 12 layers with a hidden size of 768. The number of encoder blocks and feature dimensions for each stage are presented in Tab. I. For the convolutional branch of SVLA, we used kernel sizes of $d_{1}=7$, $d_{2}=11$, and $d_{3}=21$. We set the default values for key hyperparameters as follows: $\alpha=0.125$, $\lambda=1.2$ and $\tau=0.1$. The remaining weights in our model were randomly initialized.
Subsequently, we employed the AdamW optimizer with a weight decay of 0.01. The learning rate was initialized to 3e-5 and scheduled using polynomial decay with a power of 0.9. All models were trained for 100 epochs with a batch size of 16. Images were uniformly resized to $480\times 480$ before inputting into the model, with no additional data augmentation applied.
IV-D Comparison with the State-of-the-Arts
Classical Lesion Segmentation
We compared the performance of LSMS with existing segmentation models on the RefHL-Seg, QaTa-COV19, and MosMedData+ datasets, as shown in Tab. II. In scenarios where Text is not utilized, we remove the language-related components from LSMS, such as the Language-Guided Branch in SVLA and the $\mathcal{L}_{Con}$ term in the loss function. We observed that LSMS outperformed all existing methods in Classical Lesion Segmentation task while minimizing computational costs. This underscores the efficiency and superiority of LSMS in understanding medical visual environment.
Text-Augmented Lesion Segmentation
The QaTa-COV19 and MosMedData+ datasets have been extended with textual annotations for training and validating Text-Augmented Lesion Segmentation. In Tab. II, our LSMS achieves state-of-the-art (SOTA) performance. As the proportion of textual input increases, LSMS consistently improves across all evaluation metrics. This indicates that LSMS, with textual assistance, can complement visual information with linguistic information, leading to a more accurate segmentation of the lesions.
Referring Lesion Segmentation
The RefHL-Seg dataset we constructed serves the specific purpose of training and validating the RLS task. Each sample in the dataset necessitates the model to segment the particular lesion indicated by the reference expression, thus rendering the performance of all methods suboptimal when text input is absent. Our LSMS serves as a novel baseline for the RLS task, demonstrating its significant advantages over methods applied to previous related tasks. To provide a comprehensive comparison, we evaluate LSMS against existing vision-language models on the newly proposed task and dataset. When provided with complete inputs containing both text and images, our LSMS significantly improves accuracy compared to all existing models. Specifically, LSMS achieves an increase of 8.29% over LViT [5] and 6.72% over RecLMIS [8] in mIoU, highlighting its effectiveness and superior performance in this new benchmark. This improvement stems from LSMS’s enhanced understanding of the complex medical visual environment and its precise localization capability based on linguistic cues. In Tab. II, LSMS (1/4) and LSMS (1/2) denote the models tested with only 25% and 50% of the textual input, respectively, by omitting portions such as size or shape descriptions. The results indicate that LSMS still outperforms existing methods significantly. This underscores that our LSMS, trained to precisely locate specified lesions in images with minimal textual guidance, exhibits outstanding language-guided lesion localization capability.
Figure 6: Visualization of the feature maps from different stages in LSMS. The red regions denote the Ground Truth, while the green regions represent the segmentation results of our LSMS. In sample (b), for ease of observation, the key region within the image have been enlarged, with a white rectangular box serving as a reference for location.
IV-E Ablation Study
Kernel size of the single ConvUnit in SVLA
We have meticulously evaluated the performance of using different convolutional kernels in SVLA on the validation set of RefHL-Seg dataset, as illustrated in Tab. III(a). $d_{j}=N$ indicates the ConvUnit containing a $1\times N$ convolution and a $N\times 1$ convolution. The strip-like convolution kernels aim to obtain detailed local visual information with low costs. We employed a single ConvUnit in SVLA to evaluate the impact of various convolution kernel sizes $d_{j}$. In Tab. III(a), sizes 7 and 21 show superior performance among those with comparable computational costs. The performance of medium-size convolution kernels is similar, so we choose size 11 due to its lower computational cost. The utilization of diverse convolution kernels can aid in capturing features from varying receptive fields, which helps in extracting rich local visual features. Therefore, we selected kernel sizes of $d_{1}=7$, $d_{2}=11$, and $d_{3}=21$ as the default settings.
Ablation on SVLA design
[…]
Effectiveness of FSD
To assess the effectiveness of the FSD, we compared the performance of the complete LSMS with LSMS (w/o FSD), and the results are reported in Tab. III(c). Experimental findings indicate that removing the FSD module leads to performance degradation, with Dice decrease of 3.42% and mIoU decrease of 4.54%, respectively. Furthermore, FSD incurs only a modest increase of 2.1M parameters and 1.7G Flops, yet significantly enhances performance. This underscores its efficiency in boosting performance by globally modeling multi-modal information.
Ablation on FSD design
To validate the effectiveness of the individual components in the design of FSD, we conducted ablation experiments on its constituents, as presented in Tab. III(d). It is evident that relying solely on features from the final stage is insufficient, and while adding MLP layers before concatenation increases complexity, it results in the loss of valuable information from each stage. The incorporation of the Hamburger Head design enhances the ability to globally model multi-scale information, consequently improving segmentation performance. The $P_{1},P_{2},P_{3},P_{4}$-Concat-Head-MLP design demonstrates the best performance.
FSD on various stages
Given the multi-modal features from different stages after Position Alignment, FSD concatenates them and performs joint refinement in a single forward pass. Here, $P_{i}$, $i\in\{1,2,3,4\}$, denotes the multi-modal features input to FSD from the $i$-th stage. Tab. III(e) compares multiple input sequences, confirming the value of multi-scale interaction for global reasoning. As shown, $P_{2},P_{3},P_{4}$ exhibit optimal performance for FSD. This superiority is attributed to the insufficient depth of interaction between visual information and linguistic cues in the shallow feature $P_{1}$, which include irrelevant information and hindering the localization and segmentation of specified lesions. Conversely, the latter three layers demonstrate strong visual-linguistic consistency, facilitating favorable predictions. Visualization analysis of each feature map is detailed in Sec. IV.F.
IV-F Visualization Analysis
In Fig. 6, we illustrate segmentation results and feature maps obtained from two pairs of inputs, denoted as (a) and (b). In Fig. 6(a), the language expression is “An irregular mass in the left lobe of the liver, with pronounced inhomogeneity.” The language expression for Fig. 6(b) is “The smaller low-density nodule, 27 mm in diameter.” The labels $P_{i},i\in\{1,2,3,4\}$ represent encoded multi-modal features from different stages. The segmentation results demonstrate LSMS’s ability to accurately locate and segment the specified lesions based on the language expressions. Analyzing $P_{i},i\in\{1,2,3,4\}$ reveals LSMS’s progressive focus from shallow to deep layers onto the corresponding lesions: $P_{1}$ comprehends the overall visual context, $P_{2}$ extensively attends to various objects within the image, $P_{3}$ narrows down to candidate lesions, and $P_{4}$ precisely focuses on the lesion specified by the language input. In sample (a), where only one large lesion is present, $P_{3}$ rapidly focuses on the target area. As modeling deepens, $P_{4}$ demonstrates profound visual-linguistic cues. In sample (b), where multiple lesions coexist in the image, $P_{3}$ exhibits multiple attention points. Through further vision-language interaction, $P_{4}$ can focus on the lesions specified by the expression, aiding the model in making correct predictions.
V Conclusion
In this paper, we introduce a new task, Referring Lesion Segmentation, driven by clinical demands. To support this task, we develop the RefHL-Seg benchmark for training and validating RLS. We propose LSMS for lesion segmentation with two appealing designs. LSMS integrates a novel attention mechanism to enhance object localization by tightly interacting scale-aware visual knowledge and linguistic cues. We introduce a full-scale decoder for global modeling of multi-modal features, improving boundary prediction in segmentation. Additionally, we design a specialized loss function to enhance fine-grained discrimination. Our experimental results show LSMS outperforms existing methods in both RLS and conventional lesion segmentation tasks with lower computational costs.